# Controlling Nonequilibrium Bose-Einstein Condensation with Engineered Environments

### v3

@@ Abstract
Figure 3: (a) Asymmetry matrix for selection of five modes (blue) in a ten-site chain; transitions within $S$ are colored red. (b) Achievement of BS with requested equal occupancy (solid lines: mean-field theory, circles: quasi-exact average of 500 quantum trajectories. (a) BEC pattern vs. expectation value $\langle\hat{\sigma}_{T}^{z}\rangle$ of the ‘trigger’ spin for $N=10^{2}$. The grey area indicates the total value of the heat exchanged with the energy-extracting baths, normalized by its maximum.
We solve Eq. (5) for the control variables $\bm{z}$ by targeting BS into a chosen set of three modes, {|$\{\ket{0},\ket{2},\ket{3}\}$, with target condensate fractions $[\bm{\nu}_{S,0},\bm{\nu}_{S,2},\bm{\nu}_{S,3}]=[1/10,3/10,6/10]$, finding the control values $\chi_{\ell}$ and the asymmetry matrix depicted in Fig. 2(b). The latter features the loop structure of Fig. 1(b)-top within the set of selected states (red arrows), needed to sustain BS. To verify the validity of the effective rates (4) derived, we compare in Fig. 2(d) the single-particle dynamics given by those rates with the master Eq. (S33), confirming good agreement. The achievement of the desired BS pattern in the steady state is shown in Fig. 2(e), where saturation of non-selected states starts for $N\gtrsim 30$ and the target fractions are reproduced faithfully already for $N=100$. The steady-state occupations are computed numerically here with different methods corresponding to the approximation layers used in designing the control protocol (recapitulated in [25]), to confirm their reliability: solution of the asymptotic large-$N$ theory of Eq. (2) underpinning the algorithm (5) via linear programming [26] (dashed lines), long-time propagation of the mean-field Eqs. (solid lines), and long-time propagation of the populations of $\hat{\rho}_{S}(t)$ from Eq. (1) via a quasi-exact Monte-Carlo-quantum-trajectory-type unravelling [21] (bullets, averaged over $10^{3}$ trajectories). Thanks to the bosonic enhancement in the many-body rates, the evolution converges rapidly to its BS steady state [Fig. 2(f)], with depletion of non-selected states taking place within few tens of tunnelling times for $N=50$ (solid lines), and faster and faster as $N$ increases [$N=500$ in Fig. 2(f)—dashed lines]. By exploring different connectivities through different energy-spacing–to–spin associations and solving for $\bm{z}$, protocols giving selection into any triplet of states, with the same occupation fractions as above, are also found [25]. Condensation in individual states can also be achieved, see Fig. 2(c) showing examples with |$\ket{1}$ and |$\ket{2}$.
@@
We assumed until now negligible particle loss, namely that the resonators’ relaxation is much slower than the engineered dissipation. In the SM [25], we analyze the impact of weak losses and the use of spin reservoirs realizing particle pumps to counteract them. If the total mean particle number is kept large, BS is to good approximation solely dictated by the matrix $A$, rather than by pump and loss, and is, thus, still successfully controlled with the above methods. This is easily explained by the fact that BS derives from terms in the mean-field equation that are quadratic in $n_{i}$, whereas pump and loss enter linearly. For the same reason, the BS steady state is attained well before substantial losses occur. Pump and loss may also be used as additional control parameters.
@@
A condensate switch. Realizing nonequilibrium phases which are per se robust, but tuned to be sensitive to few selected parameters, is a promising resource for, e.g., amplification and sensing. We can tune the bosonic system to a transition point between two (or more) BS configurations, such that the coupling to an external ‘trigger’ system determines which condensation pattern arises. For instance, in the five-site device of Fig. 2(a), if the quantum jump governed by the third spin is forced towards the ground state, the three-state BS turns into ground-state condensation. Consider then a coupling of the spin’s detuning $\delta_{3}$ to an additional trigger spin (‘$T$’), through the Hamiltonian $\hat{H}=\delta_{3}(r_{3}\hat{\sigma}_{3}^{x}+\hat{\sigma}_{3}^{z}\hat{\sigma}_{T}^{z})+\omega_{T}\hat{\sigma}_{T}^{z}/2$. Enforcing $|\mathcal{E}_{3}|\approx|E_{20}|$, the bosonic steady state then depends on the state of `$T^{\prime}$, through the renormalization of $\delta_{3}$ by $\langle\hat{\sigma}_{T}^{z}\rangle$. Depending on whether $\langle\hat{\sigma}_{T}^{z}\rangle=\pm 1$, the chain will exhibit either three-state or ground-state selection, as numerically verified in Fig. 3(c).
The system can also be thought of as a heat-transport switch exploiting the capability of multiple condensates to sustain large heat flow through the system, which owes to the transition rate between selected states being bosonically-enhanced by the macroscopic occupation of both states. Indeed, the total heat flow $\sum_{\ell\in\mathcal{X},ij}(E_{i}-E_{j})n_{i}(n_{j}+1)R^{(\ell)}_{ji}$ into the energy-extracting baths (subset $\mathcal{X}$) dramatically increases as the system transitions from one to three selected states [Fig. 3(c), gray area]. This strong sensitivity will make systems with controlled Bose-selection also highly-tuneable baths with structured spectral properties for systems coupled to them. An interesting perspective is also the classification of the nonequilibrium phase transitions among BS patterns, which may be distinguished in terms of the topology of the quantum-jump network, for instance exploiting the mathematical similarity between the BS asymmetry-matrix theory and Lotka-Volterra systems in evolutionary game theory [44].
@@ Acknowledgements.
Controlling Nonequilibrium Bose-Einstein Condensation with Engineered Environments
@@
Technische Universität Berlin, Institut für Theoretische Physik, Hardenbergstraße 36, Berlin 10623, Germany
I Mean-field and asymptotic high-density theory
For reference, we here briefly recapitulate the derivation of the mean-field rate Eqs. and of the Bose-selection (BS) conditions [Eq. (2)] used in the main text, which were derived in Refs. [20, 21]. Following Eq. (1), a dissipative Bose gas coupled weakly to a generic bath is described, in the interaction picture, by the Lindblad master equation
@@ I Mean-field and asymptotic high-density theory
where the operators $\hat{L}_{ij}=\hat{c}_{i}^{\dagger}\hat{c}_{j}$ describe a quantum jump from the single-particle eigenstate |$\ket{j}=\hat{c}_{j}^{\dagger}\ket{0}$ to |$\ket{i}=\hat{c}_{i}^{\dagger}\ket{0}$, occurring at rate $R_{ij}$. We denote with $\hat{n}_{i}=\hat{c}_{i}^{\dagger}\hat{c}_{i}$ the number operators. Equation (S6) generically describes the dynamics of the system coupled to an arbitrary environment, being it thermal or engineered, in the ultraweak-coupling limit where the full secular (rotating-wave) approximation is valid [55], namely such that the system-environment coupling is much smaller than the gaps in the system. Eq. (S6) implies a decoupling of the equations for the populations $p_{\bm{n}}=\bra{\bm{n}}\hat{\rho}(t)\ket{\bm{n}}$ in the Fock basis |$\ket{\bm{n}}$ of eigenstate occupations and for their coherences. Since the coherences fully decay, the steady states is diagonal in the Fock basis, $\hat{\rho}=\sum_{\bm{n}}p_{\bm{n}}\ket{\bm{n}}\!\bra{\bm{n}}$ |, and thus only the population evolution needs be solved to determine it. The latter follows from (S6) as [21]
$\frac{dp_{\bm{n}}(t)}{dt}=\sum_{i,j}n_{i}(n_{j}+1)\big{[}R_{ij}p_{\bm{n}_{ji}}(t)-R_{ji}p_{\bm{n}}(t)\big{]}\,$ (S7)
@@ III Numerical simulation methods
Large-N asymptotic theory. Solutions of the Bose-selection inequalities (2), which predict the steady state for large particle numbers, are solved via the linear programming method of Ref. [26]. To this end, we show that the inequalities (2), which restate here,
@@ IV Inequalities for the control variables
where the conditions $\bm{\nu}_{S}>0$ and $\bm{\nu}_{\bar{S}}=0$ have been imposed by construction through the choice of the target $\bm{\nu}$. We interpret the conditions in Eq. (IV) as constraints on $\bm{z}$, yielding the inequalities of Eq. (5), namely
@@
where $f(\bm{z})$ is a linear function of $\bm{z}$. While the goal of the linear program is to optimize the linear function $f(\bm{z})$ compatibly with the constraints, for our purposes it is sufficient to find a feasible solution satisfying the latter, and we can thus simply choose $f(\bm{z})=0$. In (IV), we have defined the matrix $\mathcal{B}$ with block structure $\mathcal{B}=[B,B]$ and the vector $\bm{n}_{\epsilon}=[\bm{0}_{\mathrm{dim}(B)},\bm{\nu}-\bm{\epsilon}]$, where $\bm{\epsilon}$ is a tolerance parameters with chosen value $\epsilon_{j}=10^{-2}$, allowing us to include the strict inequality of the second condition in (S29) as a $\leq$’ inequality in (IV). The constraint on $\bm{z}$ is imposed with $z_{\mathrm{max}}=1$ to ensure experimental practicality of the parameters found, such that the dispersive coupling does not exceed a maximum value $\chi_{\rm max}$, as discussed in the main text. The linear program of Eq. (IV) can then be solved by means of SciPy built-in routines [49].
@@ V Reservoir engineering with driven-damped spins
In this section, the use of driven-damped spins as artificial reservoirs is discussed, and the single-particle transition rates are derived. To obtain a clear picture of the approximations involved, we will highlight and number them consecutively. We consider, for simplicity, a single reservoir spin coupled dispersively to the $\ell$-th resonator in the chain. The generalization to multiple reservoirs is straightforward, since in this case their individual rates will simply sum up. The spin is driven and subject to fast decay. The Hamiltonian, in a frame rotating at the driving frequency of the qubit, reads
@@
where $\gamma_{\ell}$ is the relaxation rate of the spin and $D[\cdot]\hat{\rho}$ is a Lindblad dissipator, $D[\cdot]\hat{\rho}=(\cdot)\hat{\rho}(\cdot)^{\dagger}-(1/2)\{(\cdot)^{\dagger}(\cdot),\,\hat{\rho}\}$. By diagonalizing the spin Hamiltonian via the unitary transformation $\hat{U}_{\ell}=-\sin(\theta_{\ell})\hat{\sigma}_{\ell}^{z}+\cos(\theta_{\ell})\hat{\sigma}_{\ell}^{x}$ with $\theta_{\ell}=\arctan(r_{\ell})/2$, one obtains
@@
$\gamma_{\ell}D[\hat{U}_{\ell}^{\dagger}\hat{\sigma}_{\ell}^{-}\hat{U}_{\ell}]\hat{\rho}=\gamma_{\ell}\!\!\!\!\sum_{\begin{subarray}{c}\alpha,\beta\\ \in\{z,+,-\}\end{subarray}}\!\!\!\!K^{(\ell)}_{\alpha\beta}\big{(}\hat{\sigma}_{\ell}^{\alpha}\hat{\rho}\hat{\sigma}_{\ell}^{\beta}-\frac{1}{2}\{\hat{\sigma}_{\ell}^{\beta}\hat{\sigma}_{\ell}^{\alpha},\hat{\rho}\}\big{)}\ ,$ (S35)
@@
When the spin level splitting $\mathcal{E}_{\ell}$ is in resonance with a level spacing in the system, and the spin decay rate is larger than the effective system-spin coupling, the spin can be traced out, yielding effective transition rates between the bosonic-chain eigenstates, ruled by a Lorentzian bath spectral density. To show this, we start by formally diagonalizing the bosonic chain, $\hat{H}_{S}=\sum_{i}E_{i}\hat{c}_{i}^{\dagger}\hat{c}_{i}$ with |$\ket{i}=\hat{c}_{i}^{\dagger}\ket{v}$ the single-particle eigenstates and |$\ket{v}$ the vacuum. In the interaction picture with respect to the bare energies of the bosonic chain and spin, the Hamiltonian reads
@@
Next, we will trace out the spin, treating it as an environment degree of freedom, such that $\hat{H}$ of Eq. (S45) represents a system-bath coupling Hamiltonian. We proceed following standard Born-Markov-secular derivations [55]. The reason why this approach is justified, despite the spin being far from representing a true quantum bath, will be clarified in the following. The dynamics of the bosonic system is described by the Markovian quantum master equation for its density matrix $\hat{\rho}_{S}$ [55],
@@
where $\mathrm{tr}_{\sigma}$ and $\hat{\rho}_{\sigma}$ denote the trace over the degrees of freedom of the spin and the spin’s steady-state density matrix, respectively. $\hat{H}(t)$ represents the Hamiltonian of Eq. (S45) in the interaction picture. By adopting Eq. (S46), we are assuming the Born and Markov approximations (Approximation III). The latter are justified, if the spin state is negligibly affected by the coupling to the system, such that the states of the system (resonators) and the bath (spins) can be approximated as factorized,
$|\chi_{ij}^{(\ell)}|\ll|\mathcal{E}_{\ell}|\ .$ (S47)
Moreover, the spin must relax quickly enough to its steady state, that its state can be considered as constant over the evolution timescales of the system. These conditions imply that the decay rate of the spin, $\gamma_{\ell}$, needs to be much stronger than the effective spin-resonator coupling entering the Hamiltonian (S45),
@@
Inserting Eq. (S45) into (S46) and following standard manipulations [55], the following master equation in the Lindblad form is obtained,
$\frac{d\hat{\rho}_{S}}{dt}=R_{ij}^{(\ell)}D[\hat{c}^{\dagger}_{i}\hat{c}_{j}]\hat{\rho}_{S}+R_{ji}^{(\ell)}D[\hat{c}^{\dagger}_{j}\hat{c}_{i}]\hat{\rho}_{S}\ ,$ (S49)
@@
$R^{(\ell)}_{ji}=|\chi_{ij}^{(\ell)}|^{2}\mathcal{S}_{\ell}^{-}(E_{ij}),\quad(\mbox{for }\delta_{\ell}>0)$ (S50)
@@
To explicitly compute the spectra (S51) entering the rates, the spin’s steady state and correlation functions are needed. Consistently with the Markov approximation, we treat the spin as negligibly influenced by the coupling to the system, and compute such quantities as they are determined by free decay and coherent drive. This is achieved by solving the optical Bloch equations and using the quantum regression theorem [50, 55]. The problem is formally equivalent to that of a spontaneously decaying qubit with additional dephasing. The optical Bloch equations read

### v4

@@ Abstract
We solve Eq. (7) for the control variables $\bm{z}$ by targeting BS into a chosen set of three modes, {|$\{\ket{0},\ket{2},\ket{3}\}$, with target condensate fractions $[\bm{\nu}_{S,0},\bm{\nu}_{S,2},\bm{\nu}_{S,3}]=[1/10,3/10,6/10]$, finding the control values $\chi_{\ell}$ and the asymmetry matrix depicted in Fig. 2(b). The latter features the loop structure of Fig. 1(b)-top within the set of selected states (red arrows), needed to sustain BS. To verify the validity of the effective rates (6) derived, we compare in Fig. 2(d) the single-particle dynamics given by those rates with the master Eq. (S35), confirming good agreement. The achievement of the desired BS pattern in the steady state is shown in Fig. 2(e), where saturation of non-selected states starts for $N\gtrsim 30$ and the target fractions are reproduced faithfully already for $N=100$. The steady-state occupations are computed numerically here with different methods corresponding to the approximation layers used in designing the control protocol (recapitulated in [25]), to confirm their reliability: solution of the asymptotic large-$N$ theory of Eq. (2) underpinning the algorithm (7) via linear programming [26] (dashed lines), long-time propagation of the mean-field kinetic Eqs. (solid lines), and long-time propagation of the populations of $\hat{\rho}_{S}(t)$ from Eq. (1) via a quasi-exact Monte-Carlo-quantum-trajectory-type unravelling [21] (bullets, averaged over $10^{3}$ trajectories). Thanks to the bosonic enhancement in the many-body rates, the evolution converges rapidly to its BS steady state [Fig. 2(f)]. The depletion of non-selected states takes place within few tens of tunnelling times for $N=50$ (solid lines), and faster and faster as $N$ increases [$N=500$ in Fig. 2(f)—dashed lines]. By exploring different connectivities through different energy-spacing–to–spin associations and solving for $\bm{z}$, protocols giving selection into any triplet of states, with the same occupation fractions as above, are also found [25]. Condensation in individual states can also be achieved, see Fig. 2(c) showing examples with |$\ket{1}$ and |$\ket{2}$.
@@
Figure 3: (a) Asymmetry matrix for selection of five modes (blue) in a ten-site chain; transitions within $S$ are colored red. (b) Achievement of BS with requested equal occupancy (solid lines: mean-field theory, circles: quasi-exact average of 500 quantum trajectories. (a) BEC pattern vs. expectation value $\langle\hat{\sigma}_{T}^{z}\rangle$ of the ‘trigger’ spin for $N=10^{2}$. The grey area indicates the total value of the heat current $-\sum_{\ell\in B_{X}}j_{h}^{(\ell)}$ induced in the energy-extracting artificial baths (denoted as the set $B_{X}$), normalized by its maximum.
@@
We assumed until now negligible particle loss, namely that the resonators’ relaxation is much slower than the engineered dissipation. In the SM [25], we analyze the impact of weak losses and the use of spin reservoirs realizing particle pumps to counteract them. If the total mean particle number is kept large, BS is to good approximation solely dictated by the matrix $A$, rather than by pump and loss. It is, thus, still successfully controlled with the above methods. This is easily explained by the fact that BS derives from terms in the mean-field equation that are quadratic in $n_{i}$, whereas pump and loss enter linearly. For the same reason, the BS steady state is attained well before substantial losses occur. Pump and loss may also be used as additional control parameters.
@@
A condensate switch. Realizing nonequilibrium phases which are per se robust, but tuned to be sensitive to few selected parameters, is a promising resource for, e.g., amplification and sensing. We can use the ability to control BS to design similar conditions in the Bose condenser. The system can be tuned to a transition point between two (or more) BS configurations, such that the state of an additional “trigger” system coupled to it determines which condensation pattern arises. For instance, in the five-site device of Fig. 2(a), the three-state BS turns into ground-state condensation, if the quantum jump governed by the third reservoir-spin is forced towards the ground state. Consider then a coupling of the spin’s detuning $\delta_{3}$ to an additional trigger spin (‘$T$’), through the Hamiltonian $\hat{H}=\delta_{3}(r_{3}\hat{\sigma}_{3}^{x}+\hat{\sigma}_{3}^{z}\hat{\sigma}_{T}^{z})+\omega_{T}\hat{\sigma}_{T}^{z}/2$. Enforcing $|\mathcal{E}_{3}|\approx|E_{20}|$, the bosonic steady state then depends on the state of `$T^{\prime}$, through the renormalization of $\delta_{3}$ by $\langle\hat{\sigma}_{T}^{z}\rangle$. Depending on whether $\langle\hat{\sigma}_{T}^{z}\rangle=\pm 1$, the chain will exhibit either three-state or ground-state selection, as numerically verified in Fig. 3(c). A similar device may be used, for instance, as a switch that activates or deactivates the transport of macroscopic energy currents through the system. Indeed, the heat current $j_{h}^{(\ell)}$ generated in the system by the coupling to the $\ell$th artificial bath is given, within the mean-field theory, by $j_{h}^{(\ell)}=\mathrm{tr}[\dot{\hat{\rho}}_{S}\hat{H}_{S}]\approx\sum_{ij}(E_{j}-E_{i})n_{i}(n_{j}+1)R_{ji}^{(\ell)}$ [20]. The current is bosonically enhanced (quadratic in the $n_{i}$) for transitions between selected states, while it is only linearly enhanced for transitions between a selected and a non-selected state. When multiple condensates are present, the system has thus the ability to absorb and emit much more energy at large particle numbers. This feature may be of particular interest for the use of the BS system as a junction ruling quantum heat transport in a more complex device, or as an artificial bath with a highly-structured spectral density. The heat-current enhancement is shown in Fig. 3(c) for the five-site Bose condenser (grey shading), as the system passes from a single-condensate to a three-condensate phase. A interesting perspective in this context is also the classification of such nonequilibrium phase transitions among BS patterns. The latter may be distinguished in terms of the topology of the quantum-jump network, for instance exploiting the mathematical similarity between the BS asymmetry-matrix theory and Lotka-Volterra systems in evolutionary game theory [48].
@@ Acknowledgements.
Controlling Nonequilibrium Bose Condensation with Engineered Environments
@@
Technische Universität Berlin, Institut für Theoretische Physik, Hardenbergstrasse 36, Berlin 10623, Germany
I Brief review of the theory of Bose selection
For the manuscript to be self-contained, we here briefly recapitulate the derivation of the mean-field rate Eqs. and of the Bose-selection (BS) conditions [Eq. (2)] used in the main text, which were derived in Refs. [20, 21]. Following Eq. (1), a dissipative Bose gas coupled weakly to a generic bath is described, in the interaction picture, by the Lindblad master equation
@@ I Brief review of the theory of Bose selection
where the operators $\hat{L}_{ij}=\hat{c}_{i}^{\dagger}\hat{c}_{j}$ describe a quantum jump from the single-particle eigenstate |$\ket{j}=\hat{c}_{j}^{\dagger}\ket{v}$ to |$\ket{i}=\hat{c}_{i}^{\dagger}\ket{v}$, occurring at rate $R_{ij}$. Here, |$\ket{v}$ represents the vacuum state. For ease of notation, in this section we drop the subscript $S$ for the system density matrix $\hat{\rho}$, as compared to Eq. (1). We denote with $\hat{n}_{i}=\hat{c}_{i}^{\dagger}\hat{c}_{i}$ the number operators. Equation (S8) generically describes the dynamics of the system coupled to an arbitrary environment, being it thermal or engineered, in the ultraweak-coupling limit where the full secular (rotating-wave) approximation is valid [45], namely such that the system-environment coupling is much smaller than the gaps in the system. Eq. (S8) implies a decoupling of the equations for the populations $p_{\bm{n}}=\bra{\bm{n}}\hat{\rho}(t)\ket{\bm{n}}$ in the Fock basis |$\ket{\bm{n}}$ of eigenstate occupations and for their coherences. Since the coherences fully decay, the steady state is diagonal in the Fock basis, $\hat{\rho}_{ss}=\sum_{\bm{n}}p_{\bm{n}}\ket{\bm{n}}\!\bra{\bm{n}}$ |, and thus only the population evolution needs be solved to determine it. The latter follows from (S8) as [21]
$\frac{dp_{\bm{n}}(t)}{dt}=\sum_{i,j}n_{i}(n_{j}+1)\big{[}R_{ij}p_{\bm{n}_{ji}}(t)-R_{ji}p_{\bm{n}}(t)\big{]}\ ,$ (S9)
@@ III Numerical simulation methods
Large-N asymptotic theory. Solutions of the Bose-selection inequalities (2), which predict the steady state for large particle numbers, are solved via the linear programming method of Ref. [26]. To this end, we show that the inequalities (2), which we restate here,
@@ IV Inequalities for the control variables
where the conditions $\bm{\nu}_{S}>0$ and $\bm{\nu}_{\bar{S}}=0$ have been imposed by construction through the choice of the target $\bm{\nu}$. We interpret the conditions in Eq. (S28) as constraints on $\bm{z}$, yielding the inequalities of Eq. (7), namely
@@
where $f(\bm{z})$ is a linear function of $\bm{z}$. While the goal of the linear program is to optimize the linear function $f(\bm{z})$ compatibly with the constraints, for our purposes it is sufficient to find a feasible solution satisfying the latter, and we can thus simply choose $f(\bm{z})=0$. In (IV), we have defined the matrix $\mathcal{B}$ with block structure $\mathcal{B}=[B,B]$ and the vector $\bm{n}_{\epsilon}=[\bm{0}_{\mathrm{dim}(B)},\bm{\nu}-\bm{\epsilon}]$, where $\bm{\epsilon}$ are tolerance parameters with chosen value $\epsilon_{j}=10^{-2}$, allowing us to include the strict inequality of the second condition in (S31) as a $\leq$’ inequality in (IV). The constraint on $\bm{z}$ is imposed with $z_{\mathrm{max}}=1$ to ensure experimental practicality of the parameters found, such that the dispersive coupling does not exceed a maximum value $\chi_{\rm max}$, as discussed in the main text. The linear program of Eq. (IV) can then be solved by means of SciPy built-in routines [53].
@@ V Reservoir engineering with driven-damped spins
In this section, further details about the derivation of the rates $R_{ij}=\sum_{\ell}R_{ij}^{(\ell)}$ produced by the driven-damped spins used as artificial reservoirs are discussed. Following the main text, we highlight the approximations involved. Considering a single spin reservoir for simplicity, the Hamiltonian, in a frame rotating at the driving frequency of the spin, reads
@@
for the density matrix $\hat{\rho}$ of the coupled resonators-spins system, where $\gamma_{\ell}$ is the relaxation rate of the spin and $D[\cdot]\hat{\rho}$ is a Lindblad dissipator, $D[\cdot]\hat{\rho}=(\cdot)\hat{\rho}(\cdot)^{\dagger}-(1/2)\{(\cdot)^{\dagger}(\cdot),\,\hat{\rho}\}$. By diagonalizing the spin Hamiltonian via the unitary transformation $\hat{U}_{\ell}=-\sin(\theta_{\ell})\hat{\sigma}_{\ell}^{z}+\cos(\theta_{\ell})\hat{\sigma}_{\ell}^{x}$ with $\theta_{\ell}=\arctan(r_{\ell})/2$, one obtains
@@
$\gamma_{\ell}D[\hat{U}_{\ell}^{\dagger}\hat{\sigma}_{\ell}^{-}\hat{U}_{\ell}]\hat{\rho}=\gamma_{\ell}\!\!\!\!\sum_{\begin{subarray}{c}\alpha,\beta\\ \in\{z,+,-\}\end{subarray}}\!\!\!\!K^{(\ell)}_{\alpha\beta}\big{(}\hat{\sigma}_{\ell}^{\alpha}\hat{\rho}[\hat{\sigma}_{\ell}^{\beta}]^{\dagger}-\frac{1}{2}\{[\hat{\sigma}_{\ell}^{\beta}]^{\dagger}\hat{\sigma}_{\ell}^{\alpha},\hat{\rho}\}\big{)}\ ,$ (S37)
@@
as given in Eq. (4) of the main text. When the spin level splitting $\mathcal{E}_{\ell}$ is in resonance with a level spacing in the system, and the spin decay rate is larger than the effective system-spin coupling, the spin can be traced out, yielding effective transition rates between the bosonic-chain eigenstates, ruled by a Lorentzian bath spectral density. To show this, we start by formally diagonalizing the bosonic chain, $\hat{H}_{S}=\sum_{i}E_{i}\hat{c}_{i}^{\dagger}\hat{c}_{i}$ with |$\ket{i}=\hat{c}_{i}^{\dagger}\ket{v}$ the single-particle eigenstates and |$\ket{v}$ the vacuum. In the interaction picture with respect to the bare energies of the bosonic chain and spin, the Hamiltonian reads
@@
Next, we will trace out the spin, treating it as an environment degree of freedom, such that $\hat{H}$ of Eq. (S47) represents a system-bath coupling Hamiltonian. We proceed following standard Born-Markov-secular derivations [45], as discussed in the main text. The dynamics of the bosonic system is described by the Markovian quantum master equation for its density matrix $\hat{\rho}_{S}$ [45],
@@
where $\mathrm{tr}_{\sigma}$ and $\hat{\rho}_{\sigma}$ denote the trace over the degrees of freedom of the spin and the spin’s steady-state density matrix, respectively. $\hat{H}(t)$ represents the Hamiltonian of Eq. (S47) in the interaction picture. By adopting Eq. (S48), we are assuming the Born and Markov approximations (Approximation III). The latter are justified, if the spin state is negligibly affected by the coupling to the system, such that the states of the system (resonators) and the bath (spins) can be approximated as factorized, $|\chi_{ij}^{(\ell)}|\ll|\mathcal{E}_{\ell}|.$ Moreover, the spin must relax quickly enough to its steady state, such that its state can be considered as constant over the evolution timescales of the system. These conditions imply that the decay rate of the spin, $\gamma_{\ell}$, needs to be much stronger than the effective spin-resonator coupling entering the Hamiltonian (S47),
@@
Inserting Eq. (S47) into (S48) and following standard manipulations [45], a master equation in the Lindblad form as in Eq. (1) is obtained,
$\frac{d\hat{\rho}_{S}}{dt}=R_{ij}^{(\ell)}D[\hat{L}_{ij}]\hat{\rho}_{S}+R_{ji}^{(\ell)}D[\hat{L}^{\dagger}_{ij}]\hat{\rho}_{S}\ ,$ (S50)
@@
$R^{(\ell)}_{ji}=|\chi_{ij}^{(\ell)}|^{2}\mathcal{S}_{\ell}^{-}(E_{ij}),\quad(\mbox{for }\delta_{\ell}>0)\ .$ (S51)
@@
To explicitly compute the spectra (S52) entering the rates, the spin’s steady state and correlation functions are needed. Consistently with the Markov approximation, we treat the spin as negligibly influenced by the coupling to the system, and compute such quantities as they are determined by the free decay and the coherent drive. This is achieved by solving the optical Bloch equations and using the quantum regression theorem [47, 45]. The problem is formally equivalent to that of a spontaneously decaying qubit with additional dephasing. The optical Bloch equations read